Sherry and John Enterprises are using the kaizen approach to budgeting for 2018. The budgeted income statement for January 2018 is as follows:

        Sales (168,000 units)                          $1,080,000

        Less: Cost of goods sold                         650,000

        Gross margin                                            430,000

        Operating expenses                                 330,000

        (includes $60,000 of fixed costs)

        Operating income                                 $100,000

Under the kaizen approach, cost of goods sold and variable operating expenses are budgeted to decline by 1% per month.

What is budgeted cost of goods sold for March 2018?

$650,000

$637,065

$656,500

$643,500
Answer #1

Budgeted cost of goods sold for Feb 2018

= 650000-(650000×1%)

=$643500

Budgeted cost of goods sold for March 2018

= 643500-(643500×1%)

=$637065

Budgeted variable operating expenses for jan 2018

=330000-60000

=$270000

Budgeted variable operating expenses for March 2018

=270000×99%×99%

=$264627